### v3

# Ideal Analytic Sets

Abstract.
The aim of this paper is to give natural examples of $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete and $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete sets.
In the first part, we consider ideals on $\omega$. In particular, we show that the Hindman ideal $\mathcal{H}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete and consider a number of ideals generated in the similar fashion. Moreover, we show that the ideal $\mathcal{D}$ is also $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete.
In the second part, we focus on families of trees (on $\omega$ and $2$) containing a specific tree type. We show the connection between two topics and explore some classical tree types (like Sacks and Miller).
The work has been partially financed by grant 8211204601, MPK: 9130730000 from the Faculty of Pure and Applied Mathematics, Wrocław University of Science and Technology.
AMS Classification 2020: Primary: 03E15, 28A05, 54H05; Secondary: 03E75.
Keywords: analytic set, Borel set, Polish space, analytic-complete set, Borel reduction, ideal, Hindman ideal, tree, perfect tree, superperfect tree
1. Introduction
We will use standard set theoretic notions (mostly following [8] and [6]). In particular, $\omega$ is the first infinite cardinal, i.e. the set of all natural numbers, $\operatorname{\omega^{<\omega}}$ and $\operatorname{2^{<\omega}}2$ are sets of finite sequences of elements of $\omega$ and $\{0,1\}$, respectively. For this part assume that $\mathcal{X}$ and $\mathcal{Y}$ are Polish spaces, i.e. separable completely metrizable topological spaces. Classical examples are the real line $\mathbb{R}$, the Baire space $\omega^{\omega}$, the Cantor space $2^{\omega}$.
Definition 1.
Let $A\subseteq\mathcal{X}$, $B\subseteq\mathcal{Y}$. We say that $B$ is Borel reducible to $A$ if there is a Borel function $f:\mathcal{Y}\rightarrow\mathcal{X}$ satisfying $f^{-1}[A]=B$.
Definition 2.
We say that $A\subseteq\mathcal{X}$ is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete if $A$ is analytic and for every Polish space $\mathcal{Y}$ every analytic set $B\subseteq\mathcal{Y}$ is Borel reducible to $A$.
Fact 1.1.
If an analytic set $B$ is Borel reducible to $A$ and $B$ is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete, then $A$ is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
Note that existence of analytic non-Borel sets and closure of Borel sets under Borel maps implies that all $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete sets are not Borel. Moreover, in order to apply 1.1, we need an example of $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete set. Such an example can be found among trees over $\omega$.
Definition 3.
A set $T\subseteq\operatorname{\omega^{<\omega}}$ is a tree over $\omega$ if
$\left(\forall\sigma\in\operatorname{\omega^{<\omega}}\right)\left(\forall\tau\in\operatorname{\omega^{<\omega}}\right)\left(\sigma\in T\land\tau\subseteq\sigma\implies\tau\in T\right).$
Set of all trees over $\omega$ will be denoted by $\textrm{Tree}_{\omega}$. A body of a tree $T$ is a set
$[T]=\left\{\sigma\in\operatorname{\omega^{\omega}}:\left(\forall n\in\omega\right)\left(\sigma\upharpoonright n\in T\right)\right\}.$
Using above definition set $\textrm{Tree}_{\omega}$ can be seen as a $G_{\delta}$ subset of $P(\operatorname{\omega^{<\omega}})$. Therefore, $\textrm{Tree}_{\omega}$ is a Polish space. By $\textrm{IF}_{\omega}$ let us denote the collection of all ill-founded trees over $\omega$, i.e. all trees with non-empty body. It occurs that $\textrm{IF}_{\omega}$ is an example of $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete set we were looking for (see e.g. [8, Example 4.2.1]).
Definition 4.
We say that $A\subseteq\mathcal{X}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete if $\mathcal{X}\setminus A$ is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
In the first part of this paper we will consider ideals on natural numbers.
Definition 5.
A family $\mathcal{I}\subseteq P(\omega)$ is an ideal if $\mathcal{I}$ is closed
  • under taking subsets, i.e $(\forall I\in\mathcal{I})(\forall J\subseteq I)(J\in\mathcal{I})$,
  • under finite unions, i.e. $(\forall I,J\in\mathcal{I})(I\cup J\in\mathcal{I})$.
We say that $\mathcal{I}$ is proper if $\omega\not\in\mathcal{I}$.
Let us notice that if $\mathcal{I}$ is an ideal on $\omega$ then $\mathcal{I}\subseteq P(\omega)$. A space $P(\omega)$ is naturally (via characteristic function $A\mapsto\mathbbm{1}_{A}$) homeomorphic to the Cantor space $2^{\omega}$. In this context, questions about Borel complexity, being $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete or being $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete became natural for $\mathcal{I}$.
2. Ideals on $\omega$
In this section we will follow notation introduced in [4]. Let $\Lambda$ and $\Omega$ be countable infinite sets. Let $\mathcal{F}$ be a nonempty family of infinite subsets of $\Omega$ such that $F\backslash K\in\mathcal{F}$ for every $F\in\mathcal{F}$ and a finite set $K\subseteq\Omega$. For a function $\rho:\mathcal{F}\rightarrow[\Lambda]^{\omega}$ we define a family
$\mathcal{I}_{\rho}=\left\{A\subseteq\Lambda:\forall F\in\mathcal{F}(\rho(F)\nsubseteq A\right)\}.$
Definition 6.
Function $\rho:\mathcal{F}\rightarrow[\Lambda]^{\omega}$ is partition regular if
  • $(\forall E,F\in\mathcal{F})(E\subseteq F\implies\rho(E)\subseteq\rho(F))$
  • ${(\forall F\in\mathcal{F})(\forall A,B\in\Lambda)(\rho(F)=A\cup B}\implies\\ {(\exists E\in\mathcal{F})(\rho(E)\subseteq A\lor\rho(E)\subseteq B))}$
  • $\forall F\in\mathcal{F}\exists E\in\mathcal{F}(E\subseteq F\land\forall a\in\rho(E)\exists K\in[\Omega]^{<\omega}(a\notin\rho(E\backslash K)))$
As proven in [4, Proposition 3.3], if $\rho:\mathcal{F}\rightarrow[\Lambda]^{\omega}$ is partition regular, then family $\mathcal{I}_{\rho}$ is an ideal on $\Lambda$. Moreover, if $\mathcal{I}$ is an ideal on $\Lambda$, the function $\rho_{\mathcal{I}}:\mathcal{I}^{+}\rightarrow[\Lambda]^{\omega}$ given by
$\rho_{\mathcal{I}}(B)=B$
is partition regular and $\mathcal{I}=\mathcal{I}_{\rho_{\mathcal{I}}}$, so every ideal is representable in this way.
We will also need an option to extend $\rho$ on finite sets. Suppose $\mathcal{G}$ is a nonempty family of finite subsets of $\Omega$ and there exists a function $\lambda:\mathcal{G}\rightarrow\Lambda$ satisfying condition
(1) $(\forall F\in\mathcal{F})(\forall a\in\rho(F))(\exists G\in\mathcal{G}\cap\mathcal{P}(F))(\lambda(G)=a).$
Notice that it allows us to define $\rho^{\prime}:\mathcal{P}(\Omega)\rightarrow\mathcal{P}(\Lambda)$
$\rho^{\prime}(B)=\left\{\lambda(G):G\subseteq B\land G\in\mathcal{G}\right\},$
which is an extension of $\rho$ on entire $\mathcal{P}(\Omega)$. In considered ideals existence of function $\lambda$ (and the extension) will be clear, so we will identify $\rho$ with this extension.
2.1. Unified approach to reduction
In a few following subsections we will give examples of families, defined as above, which are $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete. In order to do that we construct reductions of $\textrm{IF}_{\omega}$. Since they are very similar to each other, it makes sense to define the reduction here and only later prove that they work for specific examples.
Fix function $\rho$ as earlier. We would like to construct a Borel function $\varphi:\textrm{Tree}_{\omega}\rightarrow\mathcal{P}(\Lambda)$ such that $\varphi^{-1}[\mathcal{I}_{\rho}^{c}]=\textrm{IF}_{\omega}$. Let us first fix an injection $\alpha:\operatorname{\omega^{<\omega}}\rightarrow\omega$ satisfying
(2) $\sigma\subseteq\tau\Rightarrow\alpha(\sigma)\leq\alpha(\tau).$
Now define
$\varphi(T)=\bigcup_{\sigma\in T}\rho(\{\alpha(\sigma\upharpoonright k):\;\;k<|\sigma|\}).$
From condition (1), if $\tau\in[T]$, then $\rho(\{\alpha(\tau\upharpoonright k):k\in\omega\})\subseteq\varphi(T)$, so image of an ill-founded tree is in $\mathcal{I}_{\rho}^{c}$. Therefore, the only thing left is to prove that if $\rho(B)\subseteq\varphi(T)$ for some $B\in\mathcal{F}$, then $[T]\neq\emptyset$.
2.2. Ramsey ideal
Definition 7.
Define function $r:[\omega]^{\omega}\rightarrow[[\omega]^{2}]^{\omega}$ by
$r(B)=[B]^{2}.$
Ideal $\mathcal{R}=\mathcal{I}_{r}$ is called the Ramsey ideal.
In [2, Theorem 3.3] the following result was already proven. However, we would like to give a proof based on the unified approach.
Theorem 2.1.
Set $\mathcal{R}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete.
Proof.
Suppose that $[B]^{2}\subseteq\varphi(T)$ for some $B\in[\omega]^{\omega}$. Take any $a,b\in B$, $a<b$. $\{a,b\}\in\varphi(T)$, hence (from definition of $\varphi$), there is $\sigma\in T$ such that $\alpha^{-1}(a)=\sigma\upharpoonright k$ and $\alpha^{-1}(b)=\sigma\upharpoonright l$ for some $k,l<|\sigma|$, $k\neq l$. Thus, because of condition (2), $\alpha^{-1}(a)\subset\alpha^{-1}(b)$ and $\alpha^{-1}(a),\alpha^{-1}(b)\in T$. Using this argument we can find a sequence $(a_{n})_{n\in\omega}$ of elements of $B$ such that
$i<j\implies\alpha^{-1}(a_{i})\subset\alpha^{-1}(a_{j})$
and $\alpha^{-1}(a_{i})\in T$ for every $i\in\omega$. That generates an infinite branch in tree $T$. ∎
2.3. Hindman ideal
Definition 8.
Define function $\textrm{FS}:[\omega]^{\omega}\rightarrow[\omega]^{\omega}$ by
$\textrm{FS}(B)=\left\{\sum_{n\in F}n:\;\;F\subseteq B\text{ is nonempty and finite}\right\}.$
A set $A\subseteq\omega$ is called an IP-set if there is an infinite $B\subseteq A$ satisfying $\textrm{FS}(B)\subseteq A$.
Hindman ideal is a family $\mathcal{H}$ of all non-IP-sets, i.e. $\mathcal{H}=\mathcal{I}_{\textrm{FS}}$. As proven in [4, Proposition 5.2.(5)], Hindman ideal is coanalytic. Below theorem is a strengthening of that result.
Theorem 2.2.
Set $\mathcal{H}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete.
Proof.
First, we need an additional condition on function $\alpha$, namely that $\alpha:\operatorname{\omega^{<\omega}}\rightarrow\left\{2^{2n}:n\in\omega\right\}$.
Suppose that $\varphi(T)$ is an IP-set, i.e. $\textrm{FS}(B)\subseteq\varphi(T)$ for some $B\in\operatorname{[\omega]^{\omega}}$. Note that $B\subseteq\textrm{FS}(B)$ and, because of the formula for $\varphi$, $\varphi(T)$ contains only numbers, which in binary expansion have $1$ at even positions. From now on we will identify natural numbers with their binary expansions. Take any $a_{0}\in B$ and consider a set $B_{0}=B\setminus\{b\in B:b>a_{0}\}$. Take any $b_{0}\in B_{0}$. From above remark $a_{0}$ and $b_{0}$ have $1$s at even positions. Moreover, positions of $1$s in $a_{0}$ are disjoint from positions of $1$s in $b_{0}$ (otherwise $a_{0}+b_{0}\in\textrm{FS}(B)$ has $1$ at an odd position). Put $a_{1}=a_{0}+b_{0}$, consider a set $B_{1}=B\setminus\{b\in B:b>a_{1}\}$ and take any $b_{1}\in B_{1}$. With the same argument as before, $b_{1}$ and $a_{1}$ have $1$s at disjoint positions. So we put $a_{2}=a_{1}+b_{1}$ and go on with the construction.
As a result we obtain a sequence $(a_{n})_{n\in\omega}$ of natural numbers satisfying
$a_{n}\text{ has 1 at position $k$}\Rightarrow a_{n+1}\text{ has 1 at position $k$}$
for every $n\in\omega$. This sequence induces a sequence $(2^{c_{n}})_{n\in\omega}$ of increasing powers of $2$ (by decomposing elements of $(a_{n})_{n\in\omega}$ into sums of powers of $2$) contained in $\varphi(T)$. Now note that, by definition of $\varphi$ and condition (2),
$\alpha^{-1}(2^{c_{n}})\in T\;\land\;\alpha^{-1}(2^{c_{n}})\subseteq\alpha^{-1}(2^{c_{n+1}})$
holds for every $n\in\omega$. Hence, $(\alpha^{-1}(2^{c_{n}}))_{n\in\omega}$ is an increasing sequence of elements of $T$, so $[T]\neq\emptyset$. ∎
2.4. Modifications of Hindman ideal
Definition 9.
For $k\geq 2$ define function $\textrm{FS}_{k}:[\omega]^{\omega}\rightarrow[\omega]^{\omega}$ by
$\textrm{FS}_{k}(B)=\left\{\sum_{n\in F}n:\;\;|F|=k\right\}.$
A set $A\subseteq\omega$ is called an $IP_{k}$-set if there is an infinite $B\subseteq A$ satisfying $\textrm{FS}_{k}(B)\subseteq A$.
From now on let us fix a natural number $k\geq 2$.
Analogously as before, $\mathcal{H}_{k}$ will denote a family of all non-$\operatorname{IP}_{k}$-sets, i.e. $\mathcal{H}_{k}=\mathcal{I}_{\textrm{FS}_{k}}$
Fact 2.3.
Family $\mathcal{H}_{k}$ forms an ideal.
Proof.
Closure under taking subsets is clear. Let $A,B\subseteq\omega$ be non-$\operatorname{IP}_{k}$-sets (without loss of generality disjoint). Suppose $A\cup B$ is an $\operatorname{IP}_{k}$-set, i.e. there is $C\in\operatorname{[\omega]^{\omega}}$ such that $\textrm{FS}_{k}(C)\subseteq A\cup B$. Consider a function $c:[C]^{k}\rightarrow\{A,B\}$ defined with formula
$c(E)=\begin{cases}A,&\mbox{if }\sum_{n\in E}n\in A\\ B,&\mbox{if }\sum_{n\in E}n\in B\\ \end{cases}.$
Using Ramsey’s Theorem, there is an infinite $D\subseteq C$ such that all $k$-element subsets of $D$ are "colored" by $c$ with the same color, without loss of generality $A$. Thus, $\textrm{FS}_{k}(D)\subseteq A$, what gives us a contradiction. ∎
Obviously, every IP-set is $\operatorname{IP}_{k}$ set, so $\mathcal{H}_{k}$ is a subideal of $\mathcal{H}$. This way we obtain an infinite collection of ideals over $\omega$. We shall first focus on the structure of this collection.
Fact 2.4.
Let $n,m\in\omega$. Then
$n\mid m\iff\mathcal{H}_{n}\subseteq\mathcal{H}_{m}.$
Proof.
Assume $n\mid m$. Take $k\in\omega$ such that $m=nk$. Suppose $A$ is an $\operatorname{IP}_{m}$-set, i.e. there is an infinite set $B\subseteq\omega$ satisfying $\textrm{FS}_{m}(B)\subseteq A$. Since $B$ is infinite, there is $p\in\omega$ such that a set
$B_{p}=\{a\in B:a\equiv p\mod n\}$
is infinite. Fix $a_{0},a_{1},\ldots,a_{n(k-1)-1}\in B_{p}$. Note that $n\mid\sum_{i<n(k-1)}a_{i}$, so $b=\frac{1}{n}\sum_{i<n(k-1)}a_{i}$ is a natural number. Now, consider sets $B^{\prime}=B\backslash\{a_{0},a_{1},\ldots,a_{n(k-1)-1}\}$ and $B^{\prime\prime}=B^{\prime}+b=\{c+b:c\in B^{\prime}\}$. Take any $b_{0},b_{1},\ldots,b_{n-1}\in B^{\prime\prime}$. There are $c_{0},c_{1},\ldots,c_{n-1}\in B^{\prime}$ such that $b_{i}=c_{i}+b$ for every $i<n$. Then
$\sum_{i<n}b_{i}=\sum_{i<n}(c_{i}+b)=bn+\sum_{i<n}c_{i}=\sum_{j<n(k-1)}a_{j}+\sum_{i<n}c_{i}\in FS_{m}(B)$
as a sum of $n(k-1)+n=nk=m$ elements from the set $B$. Therefore, $\textrm{FS}_{n}(B^{\prime\prime})\subseteq\textrm{FS}_{m}(B)\subseteq A$, so $A$ is an $\operatorname{IP}_{n}$-set.
Assume now that $n\nmid m$. Consider set $A=\{k\in\omega:n\nmid k\}$. Observe that
$\textrm{FS}_{m}(\{kn+1:k\in\omega\})\subseteq A,$
so $A$ is an $\operatorname{IP}_{m}$-set. However, it is not an $\operatorname{IP}_{n}$-set. Suppose it is. There is an infinite $B\subseteq\omega$ such that $\textrm{FS}_{n}(B)\subseteq A$. Since $B$ is infinite, we can (like before) find $a_{0},a_{1},\ldots,a_{n-1}$ such that $a_{i}\equiv a_{j}\mod n$ for each $i,j<n$. But then $\sum_{i<n}a_{i}\in FS_{n}(B)\subseteq A$ is divisible by $n$, what yields a contradiction. ∎
Corollary 2.5.
For each $k\in\omega$, $\mathcal{H}_{k}$ is a proper subideal of $\mathcal{H}$.
It is also rather easy to see, that function $\textrm{FS}_{k}$ is continuous for every $k\in\omega$ (similarly to continuity of FS, which can be found in [4, Proposition 5.2.(2)]). Hence, from [4, Proposition 5.1]), $\mathcal{H}_{k}$ is coanalytic for every $k\in\omega$.
Theorem 2.6.
Set $\mathcal{H}_{2}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete.
Proof.
Take $\alpha$ like in the proof of Theorem 2.2.
Suppose now that $\varphi(T)$ is an $\operatorname{IP}_{2}$-set, i.e. $\textrm{FS}_{2}(B)\subseteq\varphi(T)$ for some $B\in\operatorname{[\omega]^{\omega}}$. From now on, we identify natural numbers with their binary expansions and $|n|$ means length of $n$ in this expansion. Therefore, for any natural number $n$, $n=\sum_{i<|n|}n_{i}2^{i}$. Note that $\varphi(T)$ contains only numbers, which have exactly two $1$s. Moreover, these $1$s have to be at even positions.
We claim that set $B\cap\{2^{k}:k\in\omega\}$ is infinite. Suppose it is not, so without loss of generality $B$ does not contain any power of $2$. Consider a function $f_{1}(n)=\min\{i:n_{i}=1\}$. If $f_{1}[B]$ is infinite, take any $a\in B$. Choose $b\in B$ such that $f_{1}(b)>2|a|$. But then $a+b\in\textrm{FS}_{2}(B)$ and $a+b$ has more than two $1$s in binary expansion. Hence, $f_{1}[B]$ is finite, so without loss of generality $f_{1}[B]=\{k_{1}\}$. Furthermore, all elements of $B$ have the same number at position $k_{1}+1$. If this number is $1$, we get contradiction since
$(\ldots 01\underset{k_{1}}{1}\ldots)_{2}+(\ldots 01\underset{k_{1}}{1}\ldots)_{2}=(\ldots 11\underset{k_{1}}{0}\ldots)_{2},$
$(\ldots 11\underset{k_{1}}{1}\ldots)_{2}+(\ldots 11\underset{k_{1}}{1}\ldots)_{2}=(\ldots 11\underset{k_{1}}{0}\ldots)_{2},$
so sum of elements of $B$ has $1$s in adjacent positions (we can once again assume without loss of generality that all elements of $B$ have the same number at position $k_{1}+2$ and are bigger than $2^{k_{1}+3}$). Thus, all elements of $B$ have 00 at position $k_{1}+1$. By repeating above argument for $f_{2}=\min\{i>k_{1}:n_{i}=1\}$, all elements of $B$ have $1$ at some position $k_{2}>k_{1}$ and 00 at position $k_{2}+1$ and are bigger than $2^{k_{2}+3}$. But
$(1\ldots 0\underset{k_{2}}{1}\ldots 0\underset{k_{1}}{1}\ldots)_{2}+(1\ldots 0\underset{k_{2}}{1}\ldots 0\underset{k_{1}}{1}\ldots)_{2}=(1\ldots 1\underset{k_{2}}{0}\ldots 1\underset{k_{1}}{0}\ldots)_{2},$
what again leads to contradiction as elements of $B$ sum up to a number having more that two $1$s.
Define $B^{\prime}=B\cap\{2^{2k}:k\in\omega\}$. Then $\textrm{FS}_{2}(B^{\prime})\subseteq\textrm{FS}_{2}(B)\subseteq\varphi(T)$. Take any $a,b\in B^{\prime}$, $a<b$. $a+b\in\varphi(T)$, so (according to definition of $\varphi$) there are $\sigma\in T$ and $k,l<|\sigma|$ such that $\alpha^{-1}(a)=\sigma\upharpoonright k$ and $\alpha^{-1}(b)=\sigma\upharpoonright l$. Therefore (from condition (2)) $\alpha^{-1}(a)\subset\alpha^{-1}(b)$ and $\alpha^{-1}(a),\alpha^{-1}(b)\in T$. Using this argument we can find a sequence $(a_{n})_{n\in\omega}$ of elements of $B^{\prime}$ such that
$i<j\implies\alpha^{-1}(a_{i})\subset\alpha^{-1}(a_{j})$
and $\alpha^{-1}(a_{i})\in T$ for every $i\in\omega$. That generates an infinite branch in tree $T$. ∎
Conjecture 1.
Ideal $\mathcal{H}_{k}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete for every $k\geq 2$.
2.5. Differences instead of sums
Definition 10.
Define function $\Delta:[\omega]^{\omega}\rightarrow[\omega]^{\omega}$ by
$\Delta(B)=\left\{a-b:a,b\in B\land a>b\right\}.$
A set $A\subseteq\omega$ is called a D-set if there is an infinite $B\subseteq\omega$ satisfying $\Delta(B)\subseteq A$.
Let $\mathcal{D}$ denote a family of all non-D-sets. As proved in [3], $\mathcal{D}$ is a proper subideal of $\mathcal{H}$. Moreover, it is coanalytic (again due to [4, Proposition 5.2.(5)]).
Theorem 2.7.
Set $\mathcal{D}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete.
Proof.
This time we assume that $\alpha:\operatorname{\omega^{<\omega}}\rightarrow\{2^{n}:n\in\omega\}$.
Suppose now that $\Delta(B)\subseteq\varphi(T)$ for some infinite $B\subseteq\omega$. Take $n_{0}=\min\{k:(\exists n\in\omega)(2^{n}-2^{k}\in\Delta(B))\}$. Without loss of generality we can assume that $2^{n_{0}}\in B$ (otherwise if we consider $b=minB$ and $B^{\prime}=\{a+(2^{n_{0}}-b):a\in B\}$, $\Delta(B)=\Delta(B^{\prime})$). Observe that now all elements of the set $B$ are powers of $2$ (because only differences of powers of $2$ are in $\varphi(T)$). Take any $2^{n_{1}}\in B$. $2^{n_{1}}-2^{n_{0}}\in\Delta(B)$, so (from definition of $\varphi$ and condition (2)) $\alpha^{-1}(2^{n_{0}})\subseteq\alpha^{-1}(2^{n_{1}})$ and $\alpha^{-1}(2^{n_{1}})\in T$. Now take any $2^{n_{2}}\in B$. Analogically as above, $2^{n_{2}}-2^{n_{1}}\in\Delta(B)$, hence $\alpha^{-1}(2^{n_{1}})\subseteq\alpha^{-1}(2^{n_{2}})$ and $\alpha^{-1}(2^{n_{2}})\in T$.
Repeating this argument we obtain a sequence $(n_{k})_{k\in\omega}$ such that
$(\forall k\in\omega)(\alpha^{-1}(2^{n_{k}})\subseteq\alpha^{-1}(2^{n_{k+1}})\land\alpha^{-1}(2^{n_{k}})\in T).$
Therefore, $(\alpha^{-1}(2^{n_{k}}))_{k\in\omega}$ is an increasing sequence of elements of $T$, so $[T]\neq\emptyset$. ∎
3. Trees
As noted in the introduction, a tree over a countable set $A$ can be seen as a point in space $\mathcal{P}(A^{<\omega})$, which is homeomorphic to $\mathcal{P}(\omega)$. Therefore, families of trees, as subsets of this space, have similar topological structure to ideals on $\omega$. In the second part of this work we will investigate descriptive complexity of classical families of trees, like Miller or Laver. Let us start with a theorem connecting this topic with earlier results.
Theorem 3.1.
Consider function
$M(B)=\{\sigma\in B^{<\omega}:\sigma\text{ is increasing}\}.$
A family $\mathcal{I}_{M}$ is $\operatorname{\mathbf{\Pi}_{1}^{1}}$-complete.
Proof.
In fact we can prove something even stronger, namely that $\varphi^{-1}[\textrm{IF}_{\omega}]=\textrm{IF}_{\omega}$. This, supported with $\varphi[\textrm{IF}_{\omega}]\subseteq\mathcal{I}_{M}^{c}$, will give us thesis.
Suppose that $\tau\in[\varphi(T)]$. Take any $k\in\omega$. By definition of body, $\tau\upharpoonright(k+2)\in\varphi(T)$. Hence, there is $\sigma\in T$ such that $\alpha(\sigma\upharpoonright k_{1})=\tau(k)$ and $\alpha(\sigma\upharpoonright k_{2})=\tau(k+1)$ for some $k_{1},k_{2}\in\omega$. But $\tau(k+1)>\tau(k)$ (since $\tau\upharpoonright(k+2)\in\varphi(T)$ is increasing), so because of condition (2) $k_{1}<k_{2}$. Therefore, $\alpha^{-1}(\tau(k))\subset\alpha^{-1}(\tau(k+1))$, so $(\alpha^{-1}(\tau(k)))_{k\in\omega}$ is an increasing sequence of elements of $T$, so $[T]\neq\emptyset$. ∎
Clearly, $\mathcal{I}_{M}$ is not a family of trees (it does not necessarily have to be closed under taking subsequences). However, consider a function
$\psi(A)=\left\{\sigma\upharpoonright k:\sigma\in A,k<|\sigma|\right\},$
which "treefies" every subset of $\operatorname{\omega^{<\omega}}$. Let us recall that $\textrm{Tree}_{\omega}$ is a Polish space with subbase sets of the form $\{T\in\textrm{Tree}_{\omega}:\tau\in T\}$ (inherited from the space $\mathcal{P}(\operatorname{\omega^{<\omega}})$) for some fixed $\tau\in\operatorname{\omega^{<\omega}}$. Now see that
$\psi^{-1}(\{T\in\textrm{Tree}_{\omega}:\tau\in T\})=\bigcup_{\sigma\supseteq\tau}\{A\in\mathcal{P}(\operatorname{\omega^{<\omega}}):\sigma\in A\},$
so $\psi$ is continuous. Since $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-completeness is preserved by continuous maps, we can reformulate Theorem 3.1 in terms of trees.
Corollary 3.2.
The family of all trees containing $M(B)$ for some $B\in[\omega]^{\omega}$ is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
Definition 11.
We call a tree of the form $M(B)$ for some $B\in[\omega]^{\omega}$ a Mathias bush.
Moreover, if $T_{\sigma}=\{\tau\in\operatorname{\omega^{<\omega}}:\sigma\operatorname{{}^{\frown}}\tau\in T\}$ is a Mathias bush for some $\sigma\in T$, we call tree $T$ a Mathias tree.
The idea of above definitions comes from Mathias forcing. Mathias forcing consists of pairs $(s,S)$, where $s\in\operatorname{\omega^{<\omega}}$ is increasing and $S\in[\omega]^{\omega}$ such that $\max(\textrm{rng}(s))<\min(S)$. The ordering of pairs is defined as follows:
$(s,S)\leq(t,T)\iff s\subseteq t\land S\supseteq T\land\textrm{rng}(t)\backslash\textrm{rng}(s)\subseteq S\backslash T.$
Hence, if we fix a pair $(s,S)$, we can think of Mathias tree as a tree with stem $s$ and all possible $t\in\operatorname{\omega^{<\omega}}$ such that $(s,S)\leq(t,T)$ for some $T\in[\omega]^{\omega}$.
Now note that in the proof of Theorem 3.1 having stem or not does not make any difference. Therefore we can conclude the following result.
Corollary 3.3.
The family of all trees containing a Mathias tree is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
Definition 12.
A tree $T\in\textrm{Tree}_{\omega}$ is called Miller if
$(\forall\sigma\in T)(\exists\tau\supseteq\sigma)(\exists^{\infty}n\in\omega)(\tau\operatorname{{}^{\frown}}n\in T).$
A tree $T\in\textrm{Tree}_{\omega}$ is called Laver if
$(\exists\sigma\in T)(\forall\tau\supseteq\sigma)(\exists^{\infty}n\in\omega)(\tau\operatorname{{}^{\frown}}n\in T).$
Clearly, all Mathias trees are both Miller and Laver. This remark together with proof of Theorem 3.1 and earlier remarks give us
Corollary 3.4.
The family of all trees containing a Miller tree is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
The family of all trees containing a Laver tree is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
We would also like to explore some known families of trees over set $\{0,1\}$. Like before, such trees can be seen as points in space $\mathcal{P}(\operatorname{2^{<\omega}}2)$, homeomorphic to $\mathcal{P}(\omega)$ (from which we can derive a Polish space $\textrm{Tree}_{2}$). However, unlike in case of $\omega$, $\textrm{IF}_{2}$ is a $G_{\delta}$ subset of $\textrm{Tree}_{2}$.
Definition 13.
A tree $T\in\textrm{Tree}_{2}$ is called a Sacks tree (or perfect tree) if
$(\forall\sigma\in T)(\exists\tau\in T)(\tau\supseteq\sigma\land\tau\operatorname{{}^{\frown}}0\in T\land\tau\operatorname{{}^{\frown}}1\in T).$
A tree $T\in\textrm{Tree}_{2}$ is called a Silver tree if there exist a coinfinite set $A\subseteq\omega$ and a function $x:\omega\rightarrow 2$ such that
$T=\left\{\sigma\in\operatorname{2^{<\omega}}2:(\forall n\in|\sigma|\cap A)(\sigma(n)=x(n)\right\}.$
Then we call $A$ the domain of $T$ (and $\omega\backslash A$ the co-domain) and $x$ the Silver function of $T$.
Theorem 3.5.
The family $\mathcal{S}$ of all trees containing a Sacks tree is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
Proof.
For the proof we will need following lemma, which proof can be found in [8, Proposition 4.2.5].
Lemma 3.6.
Let $X$ be an uncountable Polish space. Then
$U(X)=\left\{K\in\ K(X):|K|>\omega\right\}$
is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete, where $K(X)$ is the space of all compact subsets of $X$ equipped with Vietoris topology.
Define $\varphi:K(2^{\omega})\rightarrow\textrm{Tree}_{2}$ by
$\varphi(K)=\left\{\sigma\upharpoonright k:\sigma\in K,k\in\omega\right\}.$
We need to show that $\varphi$ is Borel and $\varphi^{-1}[\mathcal{S}]=U(2^{\omega})$. Let us start with the second condition. By $\mathbb{S}a$ denote the family of all Sacks trees.
$\varphi^{-1}[\mathcal{S}]$ =$=\left\{K\in K(2^{\omega}):\varphi(K)\in\mathcal{S}\right\}$
=$=\left\{K\in K(2^{\omega}):(\exists T\in\mathbb{S}a)(T\subseteq\varphi(K))\right\}$
=$=\left\{K\in K(2^{\omega}):(\exists T\in\mathbb{S}a)([T]\subseteq[\varphi(K)])\right\}$
=$=\left\{K\in K(2^{\omega}):(\exists T\in\mathbb{S}a)([T]\subseteq K)\right\}$
=$=\left\{K\in K(2^{\omega}):|K|>\omega\right\}=U(2^{\omega}),$
where second to last equality comes from Sacks dichotomy for closed subsets of $2^{\omega}$.
Now, fix $s\in\operatorname{2^{<\omega}}2$. With $[s]_{\textrm{Tree}_{2}}$ and $[s]_{2^{\omega}}$ denote the subbase clopen sets generated by $s$ in spaces $\textrm{Tree}_{2}$ and $2^{\omega}$, respectively, i.e.
$[s]_{\textrm{Tree}_{2}}=\left\{T\in\textrm{Tree}_{2}:s\in T\right\},[s]_{2^{\omega}}=\left\{\sigma\in 2^{\omega}:\sigma\supseteq s\right\}.$
$\varphi^{-1}[[s]_{\textrm{Tree}_{2}}]$ =$=\left\{K\in K(2^{\omega}):(\exists\sigma\in K)(\exists k\in\omega)(s=\sigma\upharpoonright k)\right\}$
=$=\left\{K\in K(2^{\omega}):K\cap[s]_{2^{\omega}}\neq\emptyset\right\},$
which is an open set in Vietoris topology. Therefore, function $\varphi$ is continuous. ∎
Theorem 3.7.
The family $\mathcal{V}$ of all trees containing a Silver tree is $\operatorname{\mathbf{\Sigma}_{1}^{1}}$-complete.
Proof.
We construct a Borel function $\varphi:\textrm{Tree}_{\omega}\rightarrow\textrm{Tree}_{2}$ such that $\varphi^{-1}[\mathcal{V}]=\textrm{IF}_{\omega}$. First, define some auxiliary functions. Let $\alpha:\omega\rightarrow\operatorname{2^{<\omega}}2$ be a binary expansion of natural numbers, $\beta:\operatorname{2^{<\omega}}2\rightarrow\operatorname{2^{<\omega}}2$ be defined by formula
$\beta((a_{0}a_{1}\ldots a_{k}))=(a_{0}a_{0}a_{1}a_{1}\ldots a_{k}a_{k})$
and $\gamma:\operatorname{\omega^{<\omega}}\rightarrow\mathcal{P}(\operatorname{2^{<\omega}}2)$ be defined by
$\gamma((n_{0}n_{1}\ldots n_{k}))=\{\beta(\alpha(n_{0}))\operatorname{{}^{\frown}}01i_{0}01\operatorname{{}^{\frown}}\ldots\operatorname{{}^{\frown}}\beta(\alpha(n_{k}))\operatorname{{}^{\frown}}01i_{k}01:\\ i_{0},i_{1}\ldots,i_{k}\in\{0,1\}\}.$
Now we can define function $\varphi^{\prime}:\textrm{Tree}_{\omega}\rightarrow\mathcal{P}(\operatorname{2^{<\omega}}2)$:
$\varphi^{\prime}(T)=\bigcup_{\sigma\in T}\gamma(\sigma).$
As we can see, $\varphi^{\prime}(T)$ are not trees, because they are not closed under taking subsequences. However, we can one again fix it by "treefying" a result, i.e.
$\varphi=\psi\circ\varphi^{\prime}.$
Clearly, if $\tau=(\tau_{0}\tau_{1}\tau_{2}\ldots)\in[T]$ then $\varphi(T)$ contains a Silver tree with Silver function $x=\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}01\underline{0}01\operatorname{{}^{\frown}}\beta(\alpha(\tau_{1}))\operatorname{{}^{\frown}}01\underline{0}01\operatorname{{}^{\frown}}\ldots$ and co-domain being a set of indices of underlined 00s.
Suppose now that $\varphi(T)$ contains a Silver tree $S$. Silver function of $S$ has to be of the form $x_{S}=\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}01i_{0}01\operatorname{{}^{\frown}}\beta(\alpha(\tau_{1}))\operatorname{{}^{\frown}}01i_{1}01\operatorname{{}^{\frown}}\ldots$ for some $\tau=(\tau_{0}\tau_{1}\tau_{2}\ldots)$ and $i_{0},i_{1},\ldots\in\{0,1\}$ (because no other sequences are produced by function $\varphi$). Note that if the co-domain of $S$ is a subset of positions of $i_{k}$s in $x_{S}$, $\tau\in[T]$ (and $\tau$ can be uniquely identified from $x_{S}$). Assume then that there is such $m$ in co-domain of $S$ (denote the co-domain of $S$ as $A_{S}$). Without loss of generality $m\leq|\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}01i_{0}01|$ (argument can be repeated inside any block of the form $\beta(\alpha(\tau_{k}))\operatorname{{}^{\frown}}01i_{k}01$).
First, note that $m\neq|\beta(\alpha(\tau_{0}))|+4$ and $m\neq|\beta(\alpha(\tau_{0}))|+5$ (i.e. it can not be any of the last two positions), since function $\varphi$ fixes values at two positions after block of the form $\beta(n)\operatorname{{}^{\frown}}01i$. Consider the case $m=|\beta(\alpha(\tau_{0}))|+2$. As $m\in A_{S}$, $\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}01,\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}00\in S$. Since in Silver tree every sequence can be extended to the one, which follows Silver function, $\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}00i_{0}01\operatorname{{}^{\frown}}\beta(\alpha(\tau_{1}))\operatorname{{}^{\frown}}0\in S$. But then in $S$ there is a sequence of the form $\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}(00)^{k}\operatorname{{}^{\frown}}(11)^{l}\operatorname{{}^{\frown}}10$ (for some $k,l\in\omega$), which is not possible by the form of function $\varphi$.
Now consider the case $m=|\beta(\alpha(\tau_{0}))|+1$. Because $m\in A_{S}$, $\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}0,\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}1\in S$. But then, similarly as in previous case, $\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}11i_{0}01\beta(\alpha(\tau_{1}))0\in S$. Therefore, a sequence of the form $\beta(\alpha(\tau_{0}))\operatorname{{}^{\frown}}11\operatorname{{}^{\frown}}(00)^{k}\operatorname{{}^{\frown}}(11)^{l}\operatorname{{}^{\frown}}10$ (for some $k,l\in\omega$) is in $S$, which is again impossible.
There are two more possibilities for $m\leq|\beta(\alpha(\tau_{0}))|$: either $2\mid m$ or $2\nmid m$. Let us focus on the second one (the first one can be analyzed in the similar way). If $\beta(\alpha(\tau_{0}))$ has $1$ at the position $m$, it has $1$ also at the position $m-1$. But then sequence of the form $\beta(n)\operatorname{{}^{\frown}}10$ (for some $n\in\omega)$ is in $S$, which leads to contradiction. If $\beta(\alpha(\tau_{0}))$ has 00 at positions $m$ and $m-1$, $\beta(n)\operatorname{{}^{\frown}}00,\beta(n)\operatorname{{}^{\frown}}01\in S$ for some $n\in\omega$ and $\beta(n)\operatorname{{}^{\frown}}00\subseteq\beta(\alpha(\tau_{0}))$. There is $\upsilon\in\operatorname{2^{<\omega}}2$, consisting only of blocks $00$ and $11$, such that $\beta(\alpha(\tau_{0}))=\beta(n)\operatorname{{}^{\frown}}00\operatorname{{}^{\frown}}\upsilon$. Then $\beta(n)\operatorname{{}^{\frown}}01\operatorname{{}^{\frown}}\upsilon\operatorname{{}^{\frown}}01\in S$, so either $\beta(n)\operatorname{{}^{\frown}}0101\in S$, $\beta(n)\operatorname{{}^{\frown}}0111\in S$, $\beta(n)\operatorname{{}^{\frown}}010001\in S$ (which all lead to immediate contradiction) or $\beta(n)\operatorname{{}^{\frown}}010011\operatorname{{}^{\frown}}\mu\operatorname{{}^{\frown}}01\in S$, where $\upsilon=0011\operatorname{{}^{\frown}}\mu$. Hence, there is $a\in\omega$ such that either $1\operatorname{{}^{\frown}}\mu\operatorname{{}^{\frown}}0\subseteq\beta(a)$ or $\beta(a)\subseteq 1\operatorname{{}^{\frown}}\mu\operatorname{{}^{\frown}}0$. The first case leads to the situation that $\beta(a)$ contains block $10$ with $1$ at even position, which is not possible. However, in the second one $\beta(a)\operatorname{{}^{\frown}}10\subseteq 1\operatorname{{}^{\frown}}\mu\operatorname{{}^{\frown}}0$, which is again impossible, because then $\beta(n)\operatorname{{}^{\frown}}01001\operatorname{{}^{\frown}}\beta(a)\operatorname{{}^{\frown}}10\in S$.